# Foundations of Adaptive High-Level Tight Control of Prostate Cancer: A Path from From Terminal Disease to Chronic Condition

### v2

@@ I Introduction
The analysis that follows of our proposed adaptive high-level tight control (HLTC) treatment of resistant prosatte cancer predicts improved outcome compared to conventional therapies, either by maximizing the total drug response time before progression or, most importantly, transforming the disease into a chronic condition for a wide range of patient conditions. This prediction could serve as a guide to improve current adaptive chemotherapy treatments. We suggest that clinicians might explore the possibility of implementing HLTC, i.e. waiting for the disease to advance further rather than initiating treatment too early. This kind of “watchful waiting” approach before treatment is not uncommon in practice, at least historically [1], but we take it a large step further. Our simulations indicate that, with the approach developed in the following, 19 of the 32 patients we analyzed could potentially have become chronic prostate cancer survivors. In what follows, we present a detailed analysis of how it is possible to make the disease chronic.
@@
Within the prostate gland, prostate tissue cells are sensitive to male androgens, serving as key regulators of cell death and growth in the tissue [2]. In a healthy state, the growth and death rates of prostate tissue cells are controlled by androgen levels, ensuring glandular balance. In contrast, prostate cancer cells disregard these androgen controls and continue to grow. Hormone deprivation therapy targets androgen-dependent prostate cells by cutting off the supply of androgen growth hormones to limit their growth [3]. However, prostate cancer cells can acquire resistance to this treatment, enabling them to proliferate even in the absence of these hormones. Such resistant cells may (1) preexist within the prostate tumor prior to initial therapy due to genetic or epigenetic diversity [4, 5, 6], or (2) emerge as de novo resistant cells through natural selection [7, 8, 9, 10].
The two distinct cell types, which are sensitive (+) or resistant (-) to androgens, can compete within a shared environment, creating a scenario similar to a resource game [11]. Although the overall carrying capacity is dictated by the maximum number of cells that can occupy a particular environment, the proportion of androgen-sensitive and -resistant cells may vary on the basis of their competitive interactions.
Testosterone serves as the primary male androgen. Androgens such as testosterone attach to androgen receptors found on the surfaces of prostate cells. Through a complex signaling pathway, the occupation of these receptors stimulates the proliferation and apoptosis of androgen-sensitive cells. These cells, whose growth and demise are regulated by testosterone, are known as T+ cells due to their dependence on bound androgen receptors for growth and survival [12].
Over time, prostate tissue can develop or already contain cells (T+/-) with an increased number of androgen receptors, increasing their sensitivity to testosterone and allowing them to proliferate with lower testosterone levels. Because these cells can thrive even when testosterone is reduced, they are labeled “resistant” to the agonist at pituitary receptors drug leuprorelin [13], although their resistance is indirectly related to the drug itself [14]. This resistance can progress, leading to prostate cancer cells that can grow entirely without testosterone, known as T- cells. These cells are not effectively managed by standard androgen deprivation therapy, which means that even when administered leuprorelin levels are elevated, T- prostate cancer cells continue to proliferate.
@@
One reason patients treated with chemical castration using leuprorelin may show continuous PSA increases in metastatic castration-resistant prostate cancer (mCRPC) is the adrenal glands’ production of testosterone [15]. To combat this, abiraterone acetate is given to inhibit androgen synthesis in the adrenal glands, further lowering testosterone levels than leuprorelin alone can achieve. Abiraterone works by inhibiting the CYP17A1 enzyme, which is crucial for the production of androgen (male hormone) in the testes, adrenal glands, and prostate cancer tissues [16]. By blocking this enzyme, the drug drastically decreases androgen steroid hormones, notably testosterone, which significantly influence the growth potential of these susceptible cells.
@@
Cancer adaptive chemotherapy, as outlined by Gatenby [17], can be interpreted through the framework of Stackelberg sequential game theory [18]. In this approach, an oncologist exploits the competitive struggle for resources among cancerous cells with the goal of fostering the growth of drug-sensitive subclones. These subclones, in turn, help curb the expansion of drug-resistant populations [19, 20].
The key idea of adaptive chemotherapy is primarily the interplay between diverse cell groups, which compete for (1) a finite cell carrying capacity and (2) the degree of rivalry among cells, even when the carrying capacity has not yet been reached. In both cases, whether nearing carrying capacity or in the presence of interpopulation competition, a more strategically dominant cell subpopulation can inhibit the growth of a strategically minor subpopulation. This strategic suppression may also impede the expansion of a fast-replicating, drug-resistant variant. In scenarios characterized by competitive limitation of growth rates near carrying capacity, adaptive chemotherapy can postpone the emergence of resistant cell groups [21]. However, as we will illustrate, when competition coefficients are sufficiently in favor of sensitive cells, it is possible to continuously suppress resistant cell populations, even when they are far from reaching their carrying capacity.
In adaptive chemotherapy, the clinician dictates the treatment strategy through dose scheduling. Ideally, both dose intensity and duration could be adjusted. However, if dose intensity remains fixed, then the PSA on-level and off-level serve as two modifiable trigger parameters to modify drug dosage. Currently, in clinical adaptive treatments, these adjustable parameters have not been optimized for maximum drug efficiency or response duration [17] and are instead set at an arbitrary 50% of the baseline (initial) PSA level of the incoming patient [22].
As Hansen and Read have noted [21], it is better to select the PSA on-level for administering drugs to be above the initial baseline measurement at the beginning of treatment. This choice should ensure that the count of drug-sensitive cancer cells, T+, approaches the total carrying capacity, while T- remains significantly below it. In a straightforward model with two populations—where the first group, comprising T+ and T+/-, is sensitive to drugs, and the second group, consisting only of T-, is resistant—the PSA off-level should be nearly equal to the on-level. Both levels need to be near the carrying capacity to keep drug-sensitive cells close to their maximum and to limit the proliferation of drug-resistant cells.
These two concepts, (1) positioning the PSA trigger on level to maintain the total cell number near the total body carrying capacity and (2) setting the differential between on and off as small as possible are not new [21]. What we have further explored and propose here are:
@@ II A Simplified Stackelberg Game Theory Model of
where the PSA decay rate is $\Xi\sim 3.5\times 10^{-1}$/day (corresponding to a characteristic decay timescale of about a few days [23]) and $\lambda$ is the PSA production rate per cell, assumed uniform across sub-populations. It is crucial to note that not all cancer cells produce PSA in the same way or at consistent levels over time; here we are relying on a rather simplistic assumption for such a complex disease to guide our estimation for parameters that describe cancer dynamics. Since the rate $\Xi$ is significantly larger than the cancer cell growth rates (which ranges from $2\times 10^{-3}$ to $2\times 10^{-2}$/day, as estimated in [22]), we can further adopt the additional quasi-steady state approximation (QSA):
@@
The oncologist (the “game leader”), begins somewhere in the game at an unknown point, presented with a patient showing elevated PSA levels, that is we do NOT know the carrying capacity of the patient’s tumor. Let us consider a simplified version for the basic case of only two prostate cancer cell sub-populations under the influence of a single drug: (1) a drug-sensitive population $N_{S}(t)$, and (2) a drug-resistant population $N_{R}(t)$. In this model, we assume no treatment- or time-dependence to their phenotypes. That is, there is no de novo resistance evolution: no phenotype switching or mutagentic response in the cancer cells to hormone deprivation. Both cell phenotypes share a common cell carrying capacity $K$ in the absence of treatment. The cancer dynamics is then described by the following two simple coupled ODEs [22]:
@@
in which $r_{S},r_{R}>0$ are maximum growth rates of sensitive and resistant cells; the cells compete for a resource (testosterone) with the competition coefficient $\alpha_{RS}\geq 0$. This coefficient represent how the two cell types interact and affect each other indirectly through their distinct metabolisms and influence on the tumor microenvironment [24]. In Eq. (3), the effect of the drug is to reduce the carrying capacity for sensitive cells by a factor of $\gamma\Lambda(t)$, where $\Lambda(t)$ denotes the time-dependent level of the drug administered to the patient. The value $\gamma$ represents the drug effectiveness, $0\leq\gamma\leq 1$. We take $\Lambda(t)$ to be binary: $\Lambda(t)=0$ during periods of no drug administration, and $\Lambda(t)=1$ during periods of (maximum) dosage. The use of $\gamma\Lambda(t)$ may appear ad-hoc, but it is predicated the presence of Abiraterone, which reduces the carrying capacity by a factor of $(1-\gamma)$ for only drug-sensitive T+ populations, but stays unchanged for resistant T- populations.
@@
Eq. (2) and Eq. (3) contain six parameters, i.e. $\{r_{S},r_{R},\alpha_{RS},\gamma,K,\lambda\}$. Presumably all these parameters are patient-specific. In addition, there are at least two more unknowns, related to the initial cancer population sizes $\{x_{S}(0),x_{R}(0)\}$ for each patient. It is important to note that at $t=0$ for each patient, they have already received other treatments (e.g., chemical castration) in preparation for the study. The modeler must determine parameter values of each patient. In previous work by Zhang et al. [25], finding unique patient-specific parameters was not done systematically. Simulations were carried out assuming all patients entered the trial at roughly half of that value, guaranteeing that adaptive therapy would have a profound effect. This assumption of being close to carrying capacity in the simulations was carried forward into the clinical trials, where PSA levels were set to a normalized value of 1 for all patients to be near the carrying capacity. There was no attempt to actually determine carrying capacity from a dose-response experiment. In subsequent work by Zhang et al. [22], the parameters have been determined systematically but under an arbitrary choice for drug effectiveness and an assumption that the cancer growth rates are identical across patients, which is not supported by the data. Furthermore, the trigger points for on/off for abiraterone therapy were arbitrarily set (the PSA off-level is half that of the PSA on-level), with no attempt to optimize these values.
@@
Hansen and Read et al. [21] realized that the Zhang et al. [22] lacked optimization. They pointed out that actually setting the off trigger close to the carrying capacity, and setting the on trig as close as possible to the one one actually could significantly improve patient time to progression, which of course makes sense since maintaining the $T^{+}$ population at as high a value as possible within the confines of this simple model can only help suppress the T- population. This was an imporyant advance, but unfortunately Hansen and Read presented no systematic procedure for actually determining carrying capacity numbers from patient data. As they point out, if one starts far from the carrying capacity and with closely matched populations of T+ and T- an adaptive protocol can actually suppress the T+ population and help the T-, a disaster.
@@ III A More Detailed Stackelberg Game Theory Model of
A central problem in achieving a truly personal, patient-specific model for treating mCRPC is the large number of free parameters. For a patient with fewer than $\sim 10$ PSA blood level measurements (typically, a patient in [22] has around $\sim 20$ measurements), the parameters in the ODEs in Eq. (2) and Eq. (3) cannot be reliably identified when fitting observational data. We therefore now consider constraining these parameters while being careful not to arbitrarily fix too many of them.
@@
$\text{PSA}_{K}\equiv\text{PSA}\Big{|}_{x=1}=\tilde{\lambda}\ .$ (6)
@@
In contrast to the previous study [22], here we acknowledge that cancer growth and patient responses to drug may vary drastically between individuals. These considerations are consistent with crude estimations of growth rates for a subset of patients [22] and with findings related to Abiraterone responses [26, 27]. We also did not fix any parameter values before fitting the model, as [22] did for $\gamma=90\%$ and $r_{S}/r_{R}=1.66$ [22]; instead, we obtain these values from the fitting procedure. Note that we did explore other options for both common parameters and patient-specific parameters, but the selection above appears to provide the best fit.
@@
The clinician goal is to prevent the patient from staying in a state of developing symptoms (e.g. as urinary issues, erectile dysfunction, pelvic pain, bone pain), which emerge when their prostate cancer have advanced far enough. We denote the PSA level at which symptoms begin to appear as $\text{PSA}_{\text{thr}}$, a value that is not known in advance. At that stage, drug administration cannot be delayed without risking further complications. For all patients in [22], Abiraterone was first administered to all patients well before any signs of symptoms appeared. We define the two policy parameters, $\eta_{\uparrow}$ and $\eta_{\downarrow}$, so that $\text{PSA}_{\uparrow}$ is a fraction $\eta_{\uparrow}$ of the clinician acceptable upper limit $\text{PSA}_{\text{thr}}$, and $\text{PSA}_{\downarrow}$ is a fraction $\eta_{\downarrow}$ of $\text{PSA}_{\uparrow}$:
@@
The values of $\eta_{\uparrow}$ and $\eta_{\downarrow}$ are therefore bounded between 00 and $1$.

### v3

# Adaptive High-Level Tight Control of Prostate Cancer: A Path from From Terminal Disease to Chronic Condition

@@ I Introduction
The analysis that follows of our proposed adaptive high-level tight control (HLTC) treatment of resistant prosatte cancer predicts improved outcome compared to conventional therapies, either by maximizing the total drug response time before progression or, most importantly, transforming the disease into a chronic condition for a wide range of patient conditions. This prediction could serve as a guide to improve current adaptive chemotherapy treatments. We suggest that clinicians might explore the possibility of implementing HLTC, i.e. waiting for the disease to advance further rather than initiating treatment too early. This kind of “watchful waiting” approach before treatment is not uncommon in practice, at least historically [chapple2002watchful], but we take it a large step further. Our simulations indicate that, with the approach developed in the following, 19 of the 32 patients we analyzed could potentially have become chronic prostate cancer survivors. In what follows, we present a detailed analysis of how it is possible to make the disease chronic.
@@
Within the prostate gland, prostate tissue cells are sensitive to male androgens, serving as key regulators of cell death and growth in the tissue [androgens]. In a healthy state, the growth and death rates of prostate tissue cells are controlled by androgen levels, ensuring glandular balance. In contrast, prostate cancer cells disregard these androgen controls and continue to grow. Hormone deprivation therapy targets androgen-dependent prostate cells by cutting off the supply of androgen growth hormones to limit their growth [deprivation]. However, prostate cancer cells can acquire resistance to this treatment, enabling them to proliferate even in the absence of these hormones. Such resistant cells may (1) preexist within the prostate tumor prior to initial therapy due to genetic or epigenetic diversity [turke2010preexistence, bhang2015studying, sottoriva2015big], or (2) emerge as de novo resistant cells through natural selection [fitzgerald2017stress, hoffmann2000environmental, hata2016tumor, ramirez2016diverse].
The two distinct cell types, which are sensitive (+) or resistant (-) to androgens, can compete within a shared environment, creating a scenario similar to a resource game [resource_game_theory]. Although the overall carrying capacity is dictated by the maximum number of cells that can occupy a particular environment, the proportion of androgen-sensitive and -resistant cells may vary on the basis of their competitive interactions.
Testosterone serves as the primary male androgen. Androgens such as testosterone attach to androgen receptors found on the surfaces of prostate cells. Through a complex signaling pathway, the occupation of these receptors stimulates the proliferation and apoptosis of androgen-sensitive cells. These cells, whose growth and demise are regulated by testosterone, are known as T+ cells due to their dependence on bound androgen receptors for growth and survival [androgen_sensitive].
Over time, prostate tissue can develop or already contain cells (T+/-) with an increased number of androgen receptors, increasing their sensitivity to testosterone and allowing them to proliferate with lower testosterone levels. Because these cells can thrive even when testosterone is reduced, they are labeled “resistant” to the agonist at pituitary receptors drug leuprorelin [leuprorelin_action], although their resistance is indirectly related to the drug itself [leuprolide_resistance]. This resistance can progress, leading to prostate cancer cells that can grow entirely without testosterone, known as T- cells. These cells are not effectively managed by standard androgen deprivation therapy, which means that even when administered leuprorelin levels are elevated, T- prostate cancer cells continue to proliferate.
@@
One reason patients treated with chemical castration using leuprorelin may show continuous PSA increases in metastatic castration-resistant prostate cancer (mCRPC) is the adrenal glands’ production of testosterone [adrenal_production]. To combat this, abiraterone acetate is given to inhibit androgen synthesis in the adrenal glands, further lowering testosterone levels than leuprorelin alone can achieve. Abiraterone works by inhibiting the CYP17A1 enzyme, which is crucial for the production of androgen (male hormone) in the testes, adrenal glands, and prostate cancer tissues [de2011abiraterone]. By blocking this enzyme, the drug drastically decreases androgen steroid hormones, notably testosterone, which significantly influence the growth potential of these susceptible cells.
@@
Cancer adaptive chemotherapy, as outlined by Gatenby [gatenby2009adaptive], can be interpreted through the framework of Stackelberg sequential game theory [stackelberg]. In this approach, an oncologist exploits the competitive struggle for resources among cancerous cells with the goal of fostering the growth of drug-sensitive subclones. These subclones, in turn, help curb the expansion of drug-resistant populations [west2020towards, west2019multidrug].
The key idea of adaptive chemotherapy is primarily the interplay between diverse cell groups, which compete for (1) a finite cell carrying capacity and (2) the degree of rivalry among cells, even when the carrying capacity has not yet been reached. In both cases, whether nearing carrying capacity or in the presence of interpopulation competition, a more strategically dominant cell subpopulation can inhibit the growth of a strategically minor subpopulation. This strategic suppression may also impede the expansion of a fast-replicating, drug-resistant variant. In scenarios characterized by competitive limitation of growth rates near carrying capacity, adaptive chemotherapy can postpone the emergence of resistant cell groups [hansen-read]. However, as we will illustrate, when competition coefficients are sufficiently in favor of sensitive cells, it is possible to continuously suppress resistant cell populations, even when they are far from reaching their carrying capacity.
In adaptive chemotherapy, the clinician dictates the treatment strategy through dose scheduling. Ideally, both dose intensity and duration could be adjusted. However, if dose intensity remains fixed, then the PSA on-level and off-level serve as two modifiable trigger parameters to modify drug dosage. Currently, in clinical adaptive treatments, these adjustable parameters have not been optimized for maximum drug efficiency or response duration [gatenby2009adaptive] and are instead set at an arbitrary 50% of the baseline (initial) PSA level of the incoming patient [zhang2022evolution].
As Hansen and Read have noted [hansen-read], it is better to select the PSA on-level for administering drugs to be above the initial baseline measurement at the beginning of treatment. This choice should ensure that the count of drug-sensitive cancer cells, T+, approaches the total carrying capacity, while T- remains significantly below it. In a straightforward model with two populations—where the first group, comprising T+ and T+/-, is sensitive to drugs, and the second group, consisting only of T-, is resistant—the PSA off-level should be nearly equal to the on-level. Both levels need to be near the carrying capacity to keep drug-sensitive cells close to their maximum and to limit the proliferation of drug-resistant cells.
These two concepts, (1) positioning the PSA trigger on level to maintain the total cell number near the total body carrying capacity and (2) setting the differential between on and off as small as possible are not new [hansen-read]. What we have further explored and propose here are:
@@ II A Simplified Stackelberg Game Theory Model of
where the PSA decay rate is $\Xi\sim 3.5\times 10^{-1}$/day (corresponding to a characteristic decay timescale of about a few days [oesterling1993effect]) and $\lambda$ is the PSA production rate per cell, assumed uniform across sub-populations. It is crucial to note that not all cancer cells produce PSA in the same way or at consistent levels over time; here we are relying on a rather simplistic assumption for such a complex disease to guide our estimation for parameters that describe cancer dynamics. Since the rate $\Xi$ is significantly larger than the cancer cell growth rates (which ranges from $2\times 10^{-3}$ to $2\times 10^{-2}$/day, as estimated in [zhang2022evolution]), we can further adopt the additional quasi-steady state approximation (QSA):
@@
The oncologist (the “game leader”), begins somewhere in the game at an unknown point, presented with a patient showing elevated PSA levels, that is we do NOT know the carrying capacity of the patient’s tumor. Let us consider a simplified version for the basic case of only two prostate cancer cell sub-populations under the influence of a single drug: (1) a drug-sensitive population $N_{S}(t)$, and (2) a drug-resistant population $N_{R}(t)$. In this model, we assume no treatment- or time-dependence to their phenotypes. That is, there is no de novo resistance evolution: no phenotype switching or mutagentic response in the cancer cells to hormone deprivation. Both cell phenotypes share a common cell carrying capacity $K$ in the absence of treatment. The cancer dynamics is then described by the following two simple coupled ODEs [zhang2022evolution]:
@@
in which $r_{S},r_{R}>0$ are maximum growth rates of sensitive and resistant cells; the cells compete for a resource (testosterone) with the competition coefficient $\alpha_{RS}\geq 0$. This coefficient represent how the two cell types interact and affect each other indirectly through their distinct metabolisms and influence on the tumor microenvironment [tumor-micro]. In Eq. (3), the effect of the drug is to reduce the carrying capacity for sensitive cells by a factor of $\gamma\Lambda(t)$, where $\Lambda(t)$ denotes the time-dependent level of the drug administered to the patient. The value $\gamma$ represents the drug effectiveness, $0\leq\gamma\leq 1$. We take $\Lambda(t)$ to be binary: $\Lambda(t)=0$ during periods of no drug administration, and $\Lambda(t)=1$ during periods of (maximum) dosage. The use of $\gamma\Lambda(t)$ may appear ad-hoc, but it is predicated the presence of Abiraterone, which reduces the carrying capacity by a factor of $(1-\gamma)$ for only drug-sensitive T+ populations, but stays unchanged for resistant T- populations.
@@
Eq. (2) and Eq. (3) contain six parameters, i.e. $\{r_{S},r_{R},\alpha_{RS},\gamma,K,\lambda\}$. Presumably all these parameters are patient-specific. In addition, there are at least two more unknowns, related to the initial cancer population sizes $\{x_{S}(0),x_{R}(0)\}$ for each patient. It is important to note that at $t=0$ for each patient, they have already received other treatments (e.g., chemical castration) in preparation for the study. The modeler must determine parameter values of each patient. In previous work by Zhang et al. [zhang_2017], finding unique patient-specific parameters was not done systematically. Simulations were carried out assuming all patients entered the trial at roughly half of that value, guaranteeing that adaptive therapy would have a profound effect. This assumption of being close to carrying capacity in the simulations was carried forward into the clinical trials, where PSA levels were set to a normalized value of 1 for all patients to be near the carrying capacity. There was no attempt to actually determine carrying capacity from a dose-response experiment. In subsequent work by Zhang et al. [zhang2022evolution], the parameters have been determined systematically but under an arbitrary choice for drug effectiveness and an assumption that the cancer growth rates are identical across patients, which is not supported by the data. Furthermore, the trigger points for on/off for abiraterone therapy were arbitrarily set (the PSA off-level is half that of the PSA on-level), with no attempt to optimize these values.
@@
Hansen and Read et al. [hansen-read] realized that the Zhang et al. [zhang2022evolution] lacked optimization. They pointed out that actually setting the off trigger close to the carrying capacity, and setting the on trig as close as possible to the one one actually could significantly improve patient time to progression, which of course makes sense since maintaining the $T^{+}$ population at as high a value as possible within the confines of this simple model can only help suppress the T- population. This was an imporyant advance, but unfortunately Hansen and Read presented no systematic procedure for actually determining carrying capacity numbers from patient data. As they point out, if one starts far from the carrying capacity and with closely matched populations of T+ and T- an adaptive protocol can actually suppress the T+ population and help the T-, a disaster.
@@ III A More Detailed Stackelberg Game Theory Model of
A central problem in achieving a truly personal, patient-specific model for treating mCRPC is the large number of free parameters. For a patient with fewer than $\sim 10$ PSA blood level measurements (typically, a patient in [zhang2022evolution] has around $\sim 20$ measurements), the parameters in the ODEs in Eq. (2) and Eq. (3) cannot be reliably identified when fitting observational data. We therefore now consider constraining these parameters while being careful not to arbitrarily fix too many of them.
@@
$\text{PSA}_{K}\equiv\text{PSA}\Big|_{x=1}=\tilde{\lambda}\$. (6)
@@
In contrast to the previous study [zhang2022evolution], here we acknowledge that cancer growth and patient responses to drug may vary drastically between individuals. These considerations are consistent with crude estimations of growth rates for a subset of patients [zhang2022evolution] and with findings related to Abiraterone responses [fizazi2017abiraterone, hashimoto2019serum]. We also did not fix any parameter values before fitting the model, as [zhang2022evolution] did for $\gamma=90\%$ and $r_{S}/r_{R}=1.66$ [zhang2022evolution]; instead, we obtain these values from the fitting procedure. Note that we did explore other options for both common parameters and patient-specific parameters, but the selection above appears to provide the best fit. [^1] For example, allowing all parameter to be patient-specific or setting $\alpha_{SR}=\alpha_{RS}=1$ do not increase the adjusted $R^{2}$-value metric in Section IV.4.
@@
The clinician goal is to prevent the patient from staying in a state of developing symptoms (e.g. as urinary issues, erectile dysfunction, pelvic pain, bone pain), which emerge when their prostate cancer have advanced far enough. We denote the PSA level at which symptoms begin to appear as $\text{PSA}_{\text{thr}}$, a value that is not known in advance. At that stage, drug administration cannot be delayed without risking further complications. For all patients in [zhang2022evolution], Abiraterone was first administered to all patients well before any signs of symptoms appeared. We define the two policy parameters, $\eta_{\uparrow}$ and $\eta_{\downarrow}$, so that $\text{PSA}_{\uparrow}$ is a fraction $\eta_{\uparrow}$ of the clinician acceptable upper limit $\text{PSA}_{\text{thr}}$, and $\text{PSA}_{\downarrow}$ is a fraction $\eta_{\downarrow}$ of $\text{PSA}_{\uparrow}$:
@@
The values of $\eta_{\uparrow}$ and $\eta_{\downarrow}$ are therefore bounded between 0 and $1$.